The ratio of two numbers is 3:4. Their sum is to the sum of their squares as 7 is to 50. Find the numbers. (Thanks for your help.)
Answer by stanbon(75887)

The ratio of two numbers is 3:4.
Let the numbers be 3x and 4x
--------------------------------------
Their sum is to the sum of their squares as 7 is to 50. Find the numbers.
Equation:
(3x+4x)/((3x)^2+(4x)^2) = 7/50
7x/25x^2 = 7/50
x/25x^2 = 1/50
1/25x = 1/50
x = 2
----------
One number is 3x = 6
The other is 4x = 8
========================
Cheers,
Stan H.